OpenStudy (anonymous):

x^4-4x^2-1=0 Solve for x

OpenStudy (kc_kennylau):

Factorize this, this is a perfect square a^2-2ab+b^2=(a-b)^2

OpenStudy (anonymous):

Put \( \large u=x^2\), then \[ \large u^2 - 4 u -1=0 \] Use the quadratic formula to find u, then find x

OpenStudy (anonymous):

\[ u=\frac{4+ \sqrt{16+4}}{2}=\frac{1}{2} \left(2 \sqrt{5}+4\right)=2+\sqrt 5>0\\ u=\frac{4- \sqrt{16+4}}{2\ }=\frac{1}{2} \left(2 \sqrt{5}-4\right)=-2+ \sqrt 5<0\\ \]

OpenStudy (anonymous):

Since u is positive, we take the first solution. Solve for x to get \[ \left\{x= -\sqrt{2+\sqrt{5}}\right\},\left\{x= \sqrt{2+\sqrt{5}}\right\} \]